OpenStudy (anonymous):

Suppose ray QS bisects angle PQR. The measure of angle PQS = (5y - 1)º and the measure of angle PQR = (8y + 12)º. Find the measure of angle PQS.

OpenStudy (anonymous):

Since we're given QS bisects PQR, we know...\[\angle{PQS}=\frac12\angle{PQR}\\\angle{PQR}=2\angle{PQS}\] Since we're also given...\[\angle{PQS}=5y-1\\\angle{PQR}=8y+12\] So, let's plug these in and solve for our variable.\[8y+12=2(5y-1)\\8y+12=10y-2\\2y=14\\y=7\]Now we can find the measure of PQR.\[\angle{PQR}=8y+12=8(7)+12=68\]

OpenStudy (anonymous):

Thanks Alot!!!! (: I had no idea what to do.

OpenStudy (anonymous):

Woops, the question asked for PQS for the final answer. \[\angle{PQS}=5(7)-1=34^\circ\]Sorry!

OpenStudy (anonymous):

Yeahh its okk.. i looked it over nd just plugged in da 7 nd got it. Thanks tho. Really (: